# Scalable Temporal Anomaly Causality Discovery in Large Systems: Achieving Computational Efficiency with Binary Anomaly Flag Data

### v2

@@ Abstract
Extracting anomaly causality facilitates diagnostics once monitoring systems detect system faults. Identifying anomaly causes in large systems involves investigating a broader set of monitoring variables across multiple subsystems. However, learning graphical causal models (GCMs) comes with a significant computational burden that restrains the applicability of most existing methods in real-time and large-scale deployments. In addition, modern monitoring applications for large systems often generate large amounts of binary alarm flags, and the distinct characteristics of binary anomaly data—the meaning of state transition and data sparsity—challenge existing causality learning mechanisms. This study proposes an anomaly causal discovery approach (AnomalyCD), addressing the accuracy and computational challenges of generating GCMs from temporal binary flag datasets. The AnomalyCD presents several strategies, such as anomaly data-aware causality testing, sparse data and prior link compression, and edge pruning adjustment approaches. We validate the performance of of the approach on two datasets: monitoring sensor data of the readout-box system of the Compact Muon Solenoid experiment at CERN, and a public data set for information technology monitoring. The results on temporal GCMs demonstrate a considerable reduction of computation overhead and a moderate enhancement of accuracy on the binary anomaly data sets. Source code: https://github.com/muleina/AnomalyCD
@@ 1 Introduction
We propose an anomaly causal discovery framework (AnomalyCD) that integrates several methods to address computational and accuracy challenges related to generating causal knowledge discovery on multivariate TS anomaly data. We employ a temporal online AD method, sparse data handling, causal graph structure learning, and causality inference using probabilistic causal modeling. Our study focuses on discovering anomaly GCM from time series (TS) binary flags, as most monitoring systems log their data in this format. Binary flag data sets are widely available, and using them for CD lessens the impact of data heterogeneity streamed from diverse AD models, providing data normalization. Although there are some recent efforts for inferring GCM from binary data [amornbunchornvej2023framework] or outlier signals [budhathoki2022causal], there is still a gap in addressing the unique challenges of binary anomaly data for large systems. Temporal binary anomaly data has distinctive characteristics different from ordinary categorical or discrete data: 1) flag transitions between normal 0 and alert $1$ have a meaning of anomaly occurrence and disappearance, and 2) data can exhibit severe sparsity because of long uniform status regimes leading to multicollinearity. The existing causality learning mechanisms have yet to be established effectively to handle these characteristics. Their accuracy suffers, and the computation cost grows quickly for large datasets, limiting their adoption in real-time applications. The AnomalyCD leverages the accuracy and alleviates the computational burden of the PCMCI algorithm by customizing the conditional independence (CI) test, and sparse data and link compression algorithms to effectively capture temporal anomaly data attributes during graph structure generation. We propose simple but novel compression algorithms for binary flag data that substantially reduce the computational cost of the GCM learning process. Moreover, we enable a Bayesian network (BN) [neapolitan2004learning], as a probabilistic causal model, to query causality inference on the TS GCM.
@@ 2.2.1 Causal Discovery in Time Series
References [runge2018causal, runge2019detecting, runge2020discovering, saggioro2020reconstructing, gerhardus2020high] introduce and extend variants of PCMCI, an extension of the PC algorithm [spirtes2000causation] leveraged with false-positive cleaning momentary conditional independence for TS CD. The PCMCI methods with linear and non-linear conditional independence tests outperform state-of-the-art techniques in causality detection on large TS data sets across a range of research fields [runge2019detecting]. Gerhardus et al. [gerhardus2020high] propose Latent-PCMCI that relaxes the causal sufficiency assumption of PC and extends PCMCI to enhance recall CD with unknown latent confounders using FCI. Saggioro et al. [saggioro2020reconstructing] extend the PCMCI to handle non-stationarity with regime-dependent GCMs using a time-windowing method. A slightly different approach is adopted in the time-aware PC (TPC) that employs the PC algorithm for TS by considering time delay, bootstrapping, and pruning [biswas2022statistical]. The approach proceeds by unrolling the TS data (adding new nodes with time delay tags) and generating a DAG by applying a set of conditions. Nauta et al. [nauta2019causal] present a temporal CD framework that consists of independent prediction convolutional networks for each variable, and attention scores are employed to infer the causal links among the variables. Training prediction models for each variable can constrain the scalability when the number of variables increases, which is the case in large complex systems.
@@ 4 Methodology
We present the AnomalyCD framework for causality diagnostics that addresses the challenges of GCM of binary anomaly data sets. The system comprises two main modules, i.e., causal graph discovery and causality inference (see Fig. 6). The causal graph discovery generates temporal causal networks and trains a BN inference model on the anomaly data streamed from AD systems. The causality inference modules handle user queries with observation conditions to provide causality and conditional probabilities using the BN model.
Figure 6: AnomalyCD: our anomaly CD framework with use-case of monitoring the HCAL systems. The online-AD detects anomalies on variables that are not actively monitored by the trained AD models of the HCAL, such as the next-generation clock and control module (ngCCM), readout module (RM), and data quality monitoring (DQM) AD models [asres2021unsupervised, asres2022long, mulugeta2022dqm, asres2024dqmtl, asres2024lightweight]
@@ 4.1 Online Anomaly Detection
where $\Lambda_{\theta}$, $\Lambda_{\iota}$ and $\Lambda_{\phi}$ denote the temporal transient, temporal trend and spectral AD flag.
@@ 5 Results and Discussion
This section presents results of the performance evaluation of our AnomalyCD and benchmark state-of-the-art methods on the HCAL-RM and EasyVista TS datasets.
@@ 5.1 Performance Metrics
We will employ qualitative approach through visual illustrations and several quantitative metrics to evaluate the proposed approaches.
@@ 5.2 Causal Discovery on the HCAL
This section will discuss the performance of the proposed AnomalyCD using a use case RBX-HEP07 system. We will first discuss the performance of the online AD approach and then proceed to the anomaly CD.
@@ 5.2.1 Online Anomaly Detection
The LHC has undergone operations that result in distinct signal patterns on the sensors (see Fig. 4). We have thus utilized change point breaks (on 2022-09-27, 2022-10-19, and 2022-11-12) in which the AD reinitializes to deal with the irrelevant data during operation interruptions.
@@
We have demonstrated the capability of our proposed online AD approach in detecting different types of anomalies on TS data with light computation overhead. The potential limitation of the approach is that it expects adequate healthy data samples, as it is challenging to detect anomalies without prior knowledge when the outlier is dominant in data that violates the rare anomaly occurrence assumption. In such cases, using trained AD models is recommended instead. Finding the optimal hyperparameters is another open challenge. Despite data normalization and standardization partially alleviating this challenge, hyperparameters tuning may require domain knowledge to capture the right data characteristics. However, hyperparameter adjustment on statistical models is lightweight and could be faster than tuning and retraining of DL models.
@@ 5.2.2 Graphical Causal Discovery
We evaluate the discovered GCM by measuring the accuracy with respect to expected interconnection graph network depicted in Fig. 10. Figure 11(a) illustrates the temporal GCM structure on the time-lag of $t=[-\tau_{\text{max}},0]$, captured by our AnomalyCD method before edge link pruning. The GCM shows expected interconnection among the sensors: the clustering of environmental temperature and humidity sensors, and the link between the temperature regulator Peltier voltage and current, and the corresponding temperature sensors. Several bidirected edges with multiple time-lags are also present in the network. Figure 11(b) shows the final temporal DAG after applying Algorithm 6 for the pruning. An interesting time-lagged causality link can be observed, where a temperature anomaly in SRT leads to anomalies in SPV and SPC; this is because the Peltier regulator increases SPV and SPC to responds to SRT’s upsurge.
@@ 5.2.3 Bayesian Causality Inference
We trained our BN model, the query engine for temporal GCM inference shown in Fig. 6, using unrolled TS data and the captured anomaly DAG skeleton. The causality queries on the BN include conditional probability (CP) inference, allowing users to query for marginal distribution of anomaly occurrence. The inference enables estimating the causal effect between two nodes given observed anomaly evidence on the others. We employ the variable elimination algorithm, an exact inference technique for solving Bayes’ equations, to estimate CPs over a subset of variables from the probabilistic graphical BN models:
@@
Table 5 presents results of anomaly causality inference queried from the $\mathcal{M}_{b}$. We quantify the causality by calculating CP of the cause or affected sensors. The CP of anomaly occurrence for the Q1T sensor increases from $0.05$ with no other evidence to above $0.90$ with the evidence of a detected anomaly flag on the related Q[2-4]T sensors at a time-lag $t=0$. The CP of an anomaly on the Q1H sensor increases from $0.26$ to above $0.85$ when there is evidence of a detected anomaly flag on the Q[2-4]H sensors at a time-lag of $t=0$. The Q1H has a higher CP with no other evidence scenario due to trend drifts. We notice the few sample differences during drift detection across the Q[1-4]H sensors lower the causality dependency strength. The CP of SPC increases to $0.45$ when SPV has an anomaly at $t=0$. The SRT is causal to the SPC at a time-lag t=1t=-1 directly and at t=5t=-5 through SPV, but with uneven edge strength of $0.32$ and $0.15$, respectively. The CP rises to $0.95$ on the SPC when anomalies are detected on both the SPV and SRT, indicating the CP is influenced by edge weight strength from multiple causal nodes.
@@ 5.3 Causal Discovery on EasyVista
Figure 14(b) presents generated the anomaly flags data using our online AD approach. We have utilized a low threshold $\alpha_{\phi}=2$ to detect more outliers beyond the main collective anomaly at indices 46,68346,68346,78346,783 to improve the CD learning. The EasyVista experts consider the PMDB and ESB variables, with highest anomaly flags, as the root causes of the anomalies [EasyRCA2023git, assaad2023root].
@@
Table 7 provide the CD performance on the raw and compressed anomaly data before and after $\mathcal{S}_{d}$, respectively. The $\mathcal{S}_{d}$ reduces the data size to $1,900$ samples ($55\%\downarrow$), which lowers computation time by 881010 times. It improves the accuracy of the causal graph by improving the precision (reducing false edges), which is demonstrated by the higher F1, and lower FPR and SHDU. The score-based HC [koller2009probabilistic] and its hybrid MMHC [tsamardinos2006max] algorithms have limited accuracy gain on the compressed data. Only the graph autoencoder (GAE) [ng2019graph] has succeeded from the DL models. The RL-BIC [zhu2019causal] and CORL [wang2021ordering] employ reinforcement learning and have missed all relevant edges. Our $\mathcal{S}_{d}$ achieves an average relative improvement of $18.31\%$, $22.05\%$, and $15.45\%$ in the F1, FPR, and SHDU, respectively, compared to raw data.
@@
Figure 16 depicts the computational cost of the GCM learning, and our AnomalyCD reduces the cost by half in memory and 10 times in processing time with $55\%$ data compression. We found the cost of the pruning $\Gamma_{e}$ negligible with less than $0.1\%$ increment in processing time.
@@ 6 Conclusion
In this study, we have presented a novel scalable framework for discovering causal graphs for binary anomaly flag datasets. The framework incorporates various approaches to tackle the computational and accuracy challenges of inferring causality in sparse binary anomaly data. We have systematically integrated the characteristics of anomaly flag data into causal condition testing, sparse data compression, prior link compression, and pruning adjustments to state-of-the-art computation efficiency and accuracy. The results of our experiments have demonstrated promising efficacy in unsupervised online anomaly detection and a significant reduction in the computational overhead associated with graphical causal discovery on time series data sets. Our approaches will facilitate real-time diagnostic tasks across various subsystems of large systems, such as the Hadron Calorimeter and other sectors related to Industry 4.0, addressing the demand for flexible, unsupervised, and lightweight approaches that are capable of handling complex and dynamic systems with limited annotated data sets. The tools presented in this study are deployed at CERN’s Large Hadron Collider to provide anomaly detection diagnostics for the calorimeter systems.
@@ Appendix A AnomalyCD Algorithms
The appendix presents our proposed algorithms for online AD approach and the anomaly CD methods that are discussed in Section 4.
@@ B.1 Causal Graph Discovery on the HCAL Dataset
The appendix presents a performance comparison of CD using different sparsity lengths $l_{m}$ on the HCAL dataset. Figure 17 illustrates the data compression and CD performance trend with $\mathcal{S}_{d}+\mathcal{S}_{e}+\Gamma_{e}$ for $l_{m}={5,10,\dots,50}$. The results demonstrate a promising and consistent gain with $\mathcal{S}_{e}$ + $\Gamma_{e}$. However, we have observed varying CD accuracy across the $l_{m}$’s, which is attributed to the dependency of the partial correlation CI test on data size. The value of $l_{m}$ impacts data compression, which leads to varying data sizes and relative frequencies of 0’s and 1’s in the binary data. The lower $l_{m}=\tau_{\text{max}}=5$ achieves the best score, as spurious links increase for longer $l_{m}$ due to higher sparsity attribution with uniform binary sequence longer than the $\tau_{\text{max}}$.

### v3

@@ Abstract
Extracting anomaly causality facilitates diagnostics once monitoring systems detect system faults. Identifying anomaly causes in large systems involves investigating a broader set of monitoring variables across multiple subsystems. However, learning graphical causal models (GCMs) comes with a significant computational burden that restrains the applicability of most existing methods in real-time and large-scale deployments. In addition, modern monitoring applications for large systems often generate large amounts of binary alarm flags, and the distinct characteristics of binary anomaly data—the meaning of state transition and data sparsity—challenge existing causality learning mechanisms. This study proposes an anomaly causal discovery approach (AnomalyCD), addressing the accuracy and computational challenges of generating GCMs from temporal binary flag datasets. The AnomalyCD presents several strategies, such as anomaly data-aware causality testing, sparse data and prior link compression, and edge pruning adjustment approaches. We validate the performance of the approach on two datasets: monitoring sensor data from the readout-box system of the Compact Muon Solenoid experiment at CERN, and a public dataset from an information technology monitoring system. The results on temporal GCMs demonstrate a considerable reduction of computation overhead and a moderate enhancement of accuracy on the binary anomaly datasets. Source code: https://github.com/muleina/AnomalyCD
@@ 1 Introduction
We propose an anomaly causal discovery approach (AnomalyCD) that integrates several methods to address computational and accuracy challenges related to generating causal knowledge discovery on multivariate TS anomaly data. We employ a temporal online AD method, sparse data handling, causal graph structure learning, and causality inference using probabilistic causal modeling. Our study focuses on discovering anomaly GCM from time series (TS) binary flags, as most monitoring systems log their data in this format. Binary flag data sets are widely available, and using them for CD lessens the impact of data heterogeneity streamed from diverse AD models, providing data normalization. Although there are some recent efforts for inferring GCM from binary data [amornbunchornvej2023framework] or outlier signals [budhathoki2022causal], there is still a gap in addressing the unique challenges of binary anomaly data for large systems. Temporal binary anomaly data has distinctive characteristics different from ordinary categorical or discrete data: 1) flag transitions between normal 0 and alert $1$ have a meaning of anomaly occurrence and disappearance, and 2) data can exhibit severe sparsity because of long uniform status regimes leading to multicollinearity. The existing causality learning mechanisms have yet to be established effectively to handle these characteristics. Their accuracy suffers, and the computation cost grows quickly for large datasets, limiting their adoption in real-time applications. The AnomalyCD leverages the accuracy and alleviates the computational burden of the PCMCI algorithm by customizing the conditional independence (CI) test, and sparse data and link compression algorithms to effectively capture temporal anomaly data attributes during graph structure generation. We propose simple but novel compression algorithms for binary flag data that substantially reduce the computational cost of the GCM learning process. Moreover, we enable a Bayesian network (BN) [neapolitan2004learning], as a probabilistic causal model, to query causality inference on the TS GCM.
@@ 2.2.1 Causal Discovery in Time Series
References [runge2018causal, runge2019detecting, runge2020discovering, saggioro2020reconstructing, gerhardus2020high] introduce and extend variants of PCMCI, an extension of the PC algorithm [spirtes2000causation] leveraged with false-positive cleaning momentary conditional independence for TS CD. The PCMCI methods with linear and non-linear conditional independence tests outperform state-of-the-art techniques in causality detection on large TS data sets across a range of research fields [runge2019detecting]. Gerhardus et al. [gerhardus2020high] propose Latent-PCMCI that relaxes the causal sufficiency assumption of PC and extends PCMCI to enhance recall CD with unknown latent confounders using FCI. Saggioro et al. [saggioro2020reconstructing] extend the PCMCI to handle non-stationarity with regime-dependent GCMs using a time-windowing method. A slightly different approach is adopted in the time-aware PC (TPC) that employs the PC algorithm for TS by considering time delay, bootstrapping, and pruning [biswas2022statistical]. The approach unrolls the TS data (adding new nodes with time delay tags) and generates a DAG by applying a set of conditions. Nauta et al. [nauta2019causal] present a temporal CD framework that consists of independent prediction convolutional networks for each variable, and attention scores are employed to infer the causal links among the variables. Training prediction models for each variable can constrain scalability when the number of variables increases, which is often the case in large complex systems.
@@ 4 Methodology
We present the AnomalyCD for causality diagnostics that addresses the challenges of GCM of binary anomaly data sets. The system comprises two main modules, i.e., causal graph discovery and causality inference (see Fig. 6). The causal graph discovery generates temporal causal networks and trains a BN inference model on the anomaly data streamed from AD systems. The causality inference modules handle user queries with observation conditions to provide causality and conditional probabilities using the BN model.
Figure 6: AnomalyCD: our anomaly CD approach with use-case of monitoring the HCAL systems. The online-AD detects anomalies on variables that are not actively monitored by the trained AD models of the HCAL, such as the next-generation clock and control module (ngCCM), readout module (RM), and data quality monitoring (DQM) AD models [asres2021unsupervised, asres2022long, mulugeta2022dqm, asres2024dqmtl, asres2024lightweight]
@@ 4.1 Online Anomaly Detection
where $\Lambda_{\theta}$, $\Lambda_{\iota}$, and $\Lambda_{\phi}$ denote the temporal transient, temporal trend, and spectral AD flags, respectively.
@@ 5 Results and Discussion
This section presents the results of the performance evaluation of our AnomalyCD and benchmark state-of-the-art methods on the HCAL-RM and EasyVista TS datasets.
@@ 5.1 Performance Metrics
We will employ visual illustrations and several quantitative metrics to evaluate the proposed approaches.
@@ 5.2 Causal Discovery on the HCAL
This section will discuss the performance of the proposed AnomalyCD using a use case HCAL-RBX system. We will first discuss the performance of the online AD approach and then proceed to the anomaly CD.
@@ 5.2.1 Online Anomaly Detection
The LHC has undergone operations that result in distinct signal patterns on the sensors (see Fig. 4). We have thus utilized change point breaks (on 2022-09-27, 2022-10-19, and 2022-11-12) in which the AD reinitializes to deal with irrelevant data during operation interruptions.
@@
We have demonstrated the capability of our proposed online AD approach in detecting different types of anomalies on TS data with light computation overhead. The potential limitation of the approach is that it expects adequate healthy data samples, as it is challenging to detect anomalies without prior knowledge when the outlier is dominant in data that violates the rare anomaly occurrence assumption. In such cases, using trained AD models is recommended. Finding the optimal hyperparameters is another open challenge. Despite data normalization and standardization partially alleviating this challenge, hyperparameters tuning may require domain knowledge to capture the right data characteristics. However, hyperparameter adjustment on statistical models is lightweight and could be faster than tuning and retraining of DL models.
@@ 5.2.2 Graphical Causal Discovery
We evaluate the discovered GCM by measuring the accuracy with respect to expected interconnection graph network depicted in Fig. 10. Figure 11(a) illustrates the temporal GCM structure on the time-lag of $t=[-\tau_{\text{max}},0]$, captured by our AnomalyCD method before edge link pruning. The GCM shows expected interconnection among the sensors: the clustering of environmental temperature and humidity sensors, and the link between the temperature regulator Peltier voltage and current, and the corresponding temperature sensors. Several bidirected edges with multiple time-lags are also present in the network. Figure 11(b) shows the final temporal DAG after applying Algorithm 6 for the pruning. An interesting time-lagged causality link can be observed, where a temperature anomaly in SRT leads to anomalies in SPV and SPC; this is because the Peltier regulator increases SPV and SPC to respond to SRT’s upsurge.
@@ 5.2.3 Bayesian Causality Inference
We trained our BN model, the query engine for temporal GCM inference shown in Fig. 6, using unrolled TS data and the captured anomaly DAG skeleton. The causality queries on the BN include conditional probability (CP) inference, allowing users to query for the marginal distribution of anomaly occurrence. The inference enables estimating the causal effect between two nodes given observed anomaly evidence on the others. We employ the variable elimination algorithm, an exact inference technique for solving Bayes’ equations, to estimate CPs over a subset of variables from the probabilistic graphical BN models:
@@
Table 5 presents results of anomaly causality inference queried from the $\mathcal{M}_{b}$. We quantify the causality by calculating the CP of the cause or affected sensors. The CP of anomaly occurrence for the Q1T sensor increases from $0.05$ with no other evidence to above $0.90$ with the evidence of a detected anomaly flag on the related Q[2-4]T sensors at a time-lag $t=0$. The CP of an anomaly on the Q1H sensor increases from $0.26$ to above $0.85$ when there is evidence of a detected anomaly flag on the Q[2-4]H sensors at a time-lag of $t=0$. The Q1H has a higher CP with no other evidence scenario due to trend drifts. We notice that the few sample differences during drift detection across the Q[1-4]H sensors lower the causality dependency strength. The CP of SPC increases to $0.45$ when SPV has an anomaly at $t=0$. The SRT is causal to the SPC at a time-lag t=1t=-1 directly and at t=5t=-5 through SPV, but with uneven edge strength of $0.32$ and $0.15$, respectively. The CP rises to $0.95$ on the SPC when anomalies are detected on both the SPV and SRT, indicating the CP is influenced by edge weight strength from multiple causal nodes.
@@ 5.3 Causal Discovery on EasyVista
Figure 14(b) presents generated the anomaly flags data using our online AD approach. We have utilized a low threshold $\alpha_{\phi}=2$ to detect more outliers beyond the main collective anomaly at indices 46,68346,68346,78346,783 to improve the CD learning. The EasyVista experts consider the PMDB and ESB variables, with the highest anomaly flags, as the root causes of the anomalies [EasyRCA2023git, assaad2023root].
@@
Table 7 provide the CD performance on the raw and compressed anomaly data before and after $\mathcal{S}_{d}$, respectively. The $\mathcal{S}_{d}$ reduces the data size to $1,900$ samples ($55\%\downarrow$), which lowers computation time by 881010 times. It increases the accuracy of the causal graph by improving the precision (reducing false edges), which is demonstrated by the higher F1, and lower FPR and SHDU. The score-based HC [koller2009probabilistic] and its hybrid MMHC [tsamardinos2006max] algorithms have limited accuracy gain on the compressed data. Only the graph autoencoder (GAE) [ng2019graph] has succeeded from the DL models. The RL-BIC [zhu2019causal] and CORL [wang2021ordering] employ reinforcement learning and have missed all relevant edges. Our $\mathcal{S}_{d}$ achieves an average relative improvement of $18.31\%$, $22.05\%$, and $15.45\%$ in the F1, FPR, and SHDU, respectively, compared to raw data.
@@
Figure 16 depicts the computational cost of the GCM learning, and our AnomalyCD reduces the memory cost by half and the processing time by 10 times with $55\%$ data compression. We found the cost of the pruning $\Gamma_{e}$ negligible with less than $0.1\%$ increment in processing time.
@@ 6 Conclusion
In this study, we have presented a novel scalable framework for discovering causal graphs from binary anomaly flag datasets. The framework incorporates various approaches to tackle the computational and accuracy challenges of inferring causality in sparse binary anomaly data. We have systematically integrated the characteristics of anomaly flag data into causal condition testing, sparse data compression, prior link compression, and pruning adjustments to achieve state-of-the-art computation efficiency and accuracy. The results of our experiments have demonstrated promising efficacy in unsupervised online anomaly detection and a significant reduction in the computational overhead associated with graphical causal discovery on time series data sets. Our approaches will facilitate real-time diagnostic tasks across various subsystems of large systems, such as the Hadron Calorimeter and other sectors related to Industry 4.0, addressing the demand for flexible, unsupervised, and lightweight approaches that are capable of handling complex and dynamic systems with limited annotated data sets. The tools presented in this study are deployed at CERN’s Large Hadron Collider to provide anomaly detection diagnostics for the calorimeter systems.
@@ Appendix A AnomalyCD Algorithms
The appendix presents our proposed algorithms for the online AD approach and the anomaly CD methods that are discussed in Section 4.
@@ B.1 Causal Graph Discovery on the HCAL Dataset
The appendix presents a performance comparison of CD using different sparsity lengths $l_{m}$ on the HCAL dataset. Figure 17 illustrates the data compression and CD performance trend with $\mathcal{S}_{d}+\mathcal{S}_{e}+\Gamma_{e}$ for $l_{m}={5,10,\dots,50}$. The results demonstrate a promising and consistent gain with $\mathcal{S}_{e}$ + $\Gamma_{e}$. However, we have observed varying CD accuracy across the $l_{m}$’s, which is attributed to the dependency of the partial correlation CI test on data size. The value of $l_{m}$ affects data compression, resulting in varying data sizes and relative frequencies of 0s and $1$s in the binary data. The lower $l_{m}=\tau_{\text{max}}=5$ achieves the best score, as spurious links increase for longer $l_{m}$ due to higher sparsity attribution with uniform binary sequences longer than the $\tau_{\text{max}}$.